# A Survey on Recent Advances in LLM-Based Multi-turn Dialogue Systems

### v1

@@ 5.1. Pipeline-based Methods
As each module in pipeline-based TOD systems is trained independently, any module’s failure in adapting to sub-tasks can result in terrible performance of the entire system. Simultaneously, as the Pipeline-based TOD system sequentially solve all the sub-tasks, errors accumulate between modules, leading to the error propagation problem. However, since each module in the Pipeline-based TOD system operates separately, ensuring consistent input and output, it becomes convenient to interchange individual modules within the Pipeline-based TOD system. With the development of PLMs, large-scale language modules fine-tuned through different approaches can be easily accessed and seamlessly integrated to replace modules within the TOD system, which enables users to easily adapt the system to sub-tasks in the target domain.
@@ 5.1.2. Dialogue State Tracking
DST methods can be divided into static ontology DST models and dynamic ontology DST models. Static ontology DST models predict dialogue state from predefined slot-value pairs and dynamic ontology DST models predict dialogue state from an unfixed set of slot-values. Many static ontology DST models (balaraman2019scalable, ; zhong2018global, ; lee2019sumbt, ) have been proposed. However, most LLM-based DST methods are based on dynamic ontology DST models, which tracks the dialogue state from unfixed slot-value pairs. For example, SAVN and MinTL (wang2020slot, ; lin2020mintl, ) focus on creating methods or frameworks where LLMs can be effectively applied. These methods achieve competitive results and allow users to plug-and-play pre-trained sequence-to-sequence models for solving DST tasks. Hu et al (hu2022context, ). proposed IC-DST, a zero-shot and few-shot DST framework based on ICL. IC-DST retrieves a few most similar turns from the labeled dialogues as prompts, which are subsequently fed into the LLMs to produce dialogue state changes of the current turn. Feng et al. proposed LDST (feng2023towards, ), a DST framework that leverage LLaMa model. LDST initially create an instruction-tuning dataset and fine-tune the LLaMa model on this dataset. Subsequently, LDST guided the LLaMa in generating accurate responses by constructing and inputting an output prompt.
@@ 5.2. End-to-End Methods
[…]
LLMs achieve significant success in ODD tasks. However, researches related to training fully LLM-based end-to-end TOD models remains relatively limited due to the lack of large amount of training datasets for TOD tasks. Therefore, most existing LLM-based end-to-end TOD methods are based on modularly end-to-end TOD systems.
@@
Peng et al. proposed Soloist (peng2021soloist, ), an approach that uses transfer learning and machine teaching to construct the end-to-end TOD system. The training process of Soloist is pretty similar to that of SimpleTOD. However, Soloist refines the data format for each dialogue turn, dialogue action $\mathcal{A}$ is no longer required. Each dialog turn in the training data can be represented as $x=[\mathcal{H},\mathcal{B},\mathcal{D},\mathcal{S}]$. The full pre-training objective of Soloist is divided into three sub-tasks: belief prediction, grounded response generation and grounded response generation. Given the length of belief state sequence $T_{\mathcal{B}}$ and the tokens before turn $t$ $\mathcal{B}_{<t}$, the objective of predicting the belief state is defined as:
@@
[…]
@@ 6. LLM Based Open-Domain Dialogue Systems
ODD systems are designed to converse on a wide range of topics without a specific task or goal. While TOD are focused on achieving specific tasks, ODD aim to provide coherent and contextually relevant responses across any topic presented by the user. ODD is primarily categorized into three methods: Retrieval-based Methods, which select responses from a predefined set; Generation-based Methods, which generate responses dynamically; and Hybrid Methods, which combine both retrieval and generation to optimize dialogue outcomes. Table 2. shows the latest advancements in these three approaches within the domain of ODD systems.
@@
[…]
@@ 6.1. Retrieval-based Methods
In ODD systems, early foundational works have set the stage for subsequent advancements. Bordes et al. (bordes2016learning, ) were shifting the focus towards end-to-end learning approaches in goal-oriented dialog systems, challenging the traditional reliance on domain-specific handcrafting. Furthering this paradigm, Tao et al. (tao2019one, ) introduced a nuanced approach to interaction depth in multi-turn response selection, demonstrating that a more profound interaction could significantly enhance context-response matching. Concurrently, Henderson et al.(henderson2017efficient, ) showcased the practical application of these concepts in a large-scale commercial environment. These seminal contributions have laid the groundwork for a new wave of innovations in retrieval-based ODD systems.
Karpukhin et al.(karpukhin2020dense, ) introduce an approach in the field of information retrieval, centering around the dense retriever model. This model represents a departure from traditional retrieval methods like Lucene-BM25, primarily through its use of dense vector representations as opposed to conventional sparse vector models such as TF-IDF and BM25. These dense representations are obtained from embeddings learned from a carefully curated set of questions and passages, leading to improvements in top-20 passage retrieval accuracy. Additionally, the model utilizes a dual-encoder framework based on BERT, with one encoder processing questions and the other focusing on passages, each mapping its input into a low-dimensional vector space for efficient retrieval.The training process of the model focuses on the optimization of embeddings to effectively align question and passage vectors. The objective function for this optimization is defined as follows:
@@
where $q_{i}$ is the question vector, $p_{i}^{+}$ the positively aligned passage vector, and $p_{i,j}^{-}$ the negative passage vectors. The similarity function $\text{sim}(q,p)$ computes the dot product of question and passage embeddings.
With the groundwork laid by dense retrieval models, subsequent efforts focused on refining the interaction depth within dialogue systems. Tao et al.(tao2019one, ) introduce the Interaction-over-Interaction (IoI) Network. IoI enhances multi-turn response selection in retrieval-based chatbots. This model deepens the interaction between context and response by leveraging multiple rounds of interaction. It introduces multiple layers of interaction between the utterance and the response, allowing the network to capture complex semantic relationships more effectively. The IoI model’s interaction blocks process the utterance-response pairs sequentially, with self-attention mechanisms and iterative refinement enhancing the depth of interaction. Additionally, the model incorporates a mechanism for aggregating matching signals across various interaction layers.
@@ 6.2. Generation-based Methods
Generation-based methods in ODD systems have evolved significantly, offering flexibility and creativity in response synthesis. Early works like Vinyals and Le (vinyals2015neural, ) and Sutskever et al. (sutskever2014sequence, ) pioneered the use of sequence-to-sequence models for coherent dialogue generation. Further advancements by Shang et al. (shang2015neural, ) introduced attention mechanisms, improving response relevance and quality, while Serban et al. (serban2016building, ) developed the Hierarchical Recurrent Encoder-Decoder for more complex dialogues. Recent progress, marked by Radford et al. (radford2019language, ), has seen the integration of large-scale transformer models like GPT-2, pushing the boundaries of fluency and contextual awareness in generated responses.
@@ 6.2.1. Knowledge-Enhanced Generation
Building on these foundational advancements in sequence-to-sequence models and attention mechanisms, Zhao et al. (zhao2020knowledge, ) propose an enhancement for dialogue generation through the integration of external knowledge with PLMs such as GPT-2. Their approach involves a BERT-based Knowledge Selection Module for selecting relevant documents $D^{\prime}$ from a set $D$ based on the dialogue context $U$. The selected knowledge is utilized in a GPT-2-based Response Generation Model for generating responses:
@@
where $g(U,D^{\prime})$ represents the integrated user dialogue context and selected knowledge, and $\theta$ denotes the GPT-2 parameters. This method allows the generation of responses that are contextually relevant and informed by external knowledge.
Moreover, while Zhao et al. focused on enhancing dialogue generation with external knowledge, Xu et al. (xu2022long, ) introduce the PLATO-LTM model, featuring a Long-Term Memory (LTM) mechanism to consistently maintain persona information. The model utilizes a Persona Extractor (PE) to classify persona labels based on user input $U_{i}$ using an ERNIE-CNN architecture. The LTM module in PLATO-LTM is responsible for matching the context with the relevant persona, enabling the system to retrieve appropriate persona-specific information. The Generation Module is represented as:
(30) $L_{\text{NLL}}=-\mathbb{E}\left[\sum_{t=1}^{T}\log p(r_{t}|c,\rho_{u},\rho_{s},r_{<t})\right].$
This equation represents the negative log-likelihood loss function, where $r_{t}$ is the generated response at time $t$, $c$ denotes the current context, $\rho_{u}$ and $\rho_{s}$ are user and system persona embeddings, respectively, and $r_{<t}$ symbolizes the responses generated up to the previous time step. PLATO-LTM’s design focuses on enhancing the coherence and engagement of conversations by dynamically managing persona information, thus facilitating more natural and contextually relevant dialogue generation.
Further expanding on the theme of personalization, Song et al. (song2020generate, ) develop the Generate-Delete-Rewrite (GDR) framework to create persona-consistent dialogues. The GDR framework operates in three stages: initially generating a response prototype, then identifying and masking elements that are inconsistent with the established persona, and finally refining the output. The process starts with the creation of an initial response vector, followed by calculating attention weights to identify inconsistencies, and concludes with a refinement stage where the initial output is adjusted to ensure persona consistency. Building on the idea of tailored dialogue generation, Tang et al. (tang2023enhancing, ) introduce the Contrastive Latent Variable (CLV)-based model. This model enhances dialogue personalization by using both sparse and dense persona resources. It begins by encoding persona information and user queries, and then generates personalized responses by combining these encoded inputs. The CLV model distinguishes itself by clustering dense persona descriptions into sparse categories, effectively utilizing varied persona data to inform and personalize dialogue responses. The CLV model is not only consistent with user personas but also enriched by a deeper understanding of individual user characteristics.

### v2

@@ 3.1. Pipeline-based Methods
As each module in pipeline-based TOD systems is trained independently, any module’s failure in adapting to sub-tasks can result in terrible performance of the entire system. Simultaneously, as the Pipeline-based TOD system sequentially solve all the sub-tasks, errors accumulate between modules, leading to the error propagation problem. However, since each module in the Pipeline-based TOD system operates separately, ensuring consistent input and output, it becomes convenient to interchange individual modules within Pipeline-based TOD systems. With the development of PLMs, large-scale modules fine-tuned through different approaches can be easily accessed and seamlessly integrated to replace modules within TOD systems, which enables users to easily adapt the system to sub-tasks in the target domain.
@@ 3.1.2. Dialogue State Tracking
DST methods can be divided into static ontology DST models and dynamic ontology DST models. Static ontology DST models predict dialogue state from predefined slot-value pairs and dynamic ontology DST models predict dialogue state from an unfixed set of slot-values. Many static ontology DST models (balaraman2019scalable, ; zhong2018global, ; lee2019sumbt, ) have been proposed.
Most LLM-based DST methods are based on dynamic ontology DST models, which tracks the dialogue state from unfixed slot-value pairs. For example, SAVN and MinTL (wang2020slot, ; lin2020mintl, ) focus on creating methods or frameworks where LLMs can be effectively applied. These methods achieve competitive results and allow users to plug-and-play pre-trained sequence-to-sequence models for solving DST tasks. Hu et al. proposed IC-DST (hu2022context, ), a zero-shot and few-shot DST framework based on ICL. IC-DST retrieves a few most similar turns from the labeled dialogues as prompts, which are subsequently fed into the LLMs to produce dialogue state changes of the current turn. Similarly, King and Flanigan proposed RefPyDST (king2023diverse, ), which re-frames DST as a Python programming task. RefPyDST searches similar in-context examples and insert them into the prompts to guide the LLMs in performing DST tasks more effectively. Inspered by these ICL-based methods, Yi et al. (yi2024intent, ) effectively integrated the output of the NLU module into dialogue history, enabling the identification of more suitable in-context examples and allowing LLMs to better extract users’ implicit intents. Feng et al. (feng2024fact, ) utilized LLMs to summarize the conversation, extracting the most critical information, and then tracked the dialogue state from the summary, thereby reducing the impact of noise in the dialogue history. Feng et al. proposed LDST (feng2023towards, ), a DST framework that leverage LLaMa model. LDST initially create an instruction-tuning dataset and fine-tune the LLaMa model on this dataset. Subsequently, LDST guided the LLaMa in generating accurate responses by constructing and inputting an output prompt.
@@ 3.2. End-to-End Methods
(11) $\mathcal{S}=\text{ End-to-end TOD }(\mathcal{H},\mathcal{KB}).$
3.2.1. Fully end-to-end TOD
@@
Peng et al. proposed Soloist (peng2021soloist, ), an approach that uses transfer learning and machine teaching to construct the end-to-end TOD system. The training process of Soloist is pretty similar to that of SimpleTOD. However, Soloist refines the data format for each dialogue turn, dialogue action $\mathcal{A}$ is no longer required. Each dialog turn in the training data can be represented as $x=[\mathcal{H},\mathcal{B},\mathcal{D},\mathcal{S}]$. The full pre-training objective of Soloist is divided into three sub-tasks: belief prediction, grounded response generation and contrastive objective. Given the length of belief state sequence $T_{\mathcal{B}}$ and the tokens before turn $t$ $\mathcal{B}_{<t}$, the objective of predicting the belief state is defined as:
@@
For the UBAR (yang2021ubar, ), previous modularly end-to-end TOD methods that are trained and evaluated in turn-level sequences where they are based on dialog history $\mathcal{H}_{t}=\left(u_{1},s_{1}\right),\left(u_{2},s_{2}\right),\ldots\left(u_{t-1},s_{t-1}\right),\left(u_{t}\right)$ to generate response in turn t. While UBAR integrates the intermediate information $\mathcal{B}$, $\mathcal{D}$, and $\mathcal{A}$ within the context. Therefore, the training sequence of UBAR in turn $t$ is de fined as $\mathcal{H}_{0},\mathcal{B}_{0},\mathcal{D}_{0},\mathcal{A}_{0},\mathcal{S}_{0},\dots\mathcal{H}_{t},\mathcal{B}_{t},\mathcal{D}_{t},\mathcal{A}_{t},\mathcal{S}_{t}]$, which is then used to fine-tune the large pre-trained model GPT-2.
Saley et al. propose SyncTOD (saley-etal-2024-synergizing, ), which aligns LLMs with the stylings of the available training data. SyncTOD generates useful hints $\hat{H}$ regarding the anticipated response. These hints enhance the quality of the examples through re-ranking and guide the LLMs (accessed via API) towards generating the expected response from within the prompt. It’s worth mentioning that aforementioned fully end-to-end TOD systems, such as SimpleTOD, train LLMs using multi-turn dialogue data, enabling them to update the dialogue state, query databases, and generate system responses step by step to complete task-oriented dialogues. However, SyncTOD concatenate the database $\mathcal{KB}$ to the LLMs prompt in JSON format. This allows the LLMs to generate system responses $\mathcal{S}$ in a single step. Experimental results indicate that this approach is not only simpler but also outperforms existing SOTA models under both few-shot and full-data settings.
King et al. (king2024unsupervised, ) introduce the first approach for building a working TOD agent with LLMs using only unlabled dialogues. Specifically, this method prompt the LLMs with a text-to-code prompt for inferring the latent dialogue state as an API call, from which the dialogue state change and dialogue acts can be derived. Then, based on this data, a multi-task fine-tuning method is utilized to train a single LLM as a complete dialogue system.
3.2.2. Modularly end-to-end TOD
@@
Nguyen el at al. propose SynTOD (chieu-etal-2024-syntod, ), a zero-shot modular end-to-end TOD system. SynTOD generates synthesis responses utilizing pre-trained LLMs, which are then used as the input to fine-tune end-to-end TOD model. Unlike most modular end-to-end TOD systems, SynTOD is composed of one encoder and two decoders. The dialogue encoder encodes the user input and dialogue history into vector representations. The belief decoder decodes the updated dialogue state and queries the database. The retrieved data, together with the vector output from the dialogue encoder, is then fed into the response decoder, enabling the model to generate dialogue actions and system responses. This procedure is repeated until the dialogue is complete.
@@ 4. LLM Based Open-Domain Dialogue Systems
ODD systems, unlike TOD systems, are designed to engage in conversations across a wide range of topics without a specific task or goal in mind. The primary aim of ODD systems is to provide coherent, contextually relevant, and natural responses, regardless of the user’s query. This is more challenging compared to TOD, as ODD must handle a broad diversity of topics and maintain context throughout the interaction.
As shown in 4, ODD systems greatly benefit from LLMs, which enable them to accurately respond to a broad spectrum of user queries, regardless of the topic. LLMs also enhance these systems’ ability to detect emotions, allowing for empathetic responses that contribute to more natural and engaging conversations. Additionally, LLMs improve personalization by tailoring recommendations based on user preferences, offering a more individualized and dynamic user experience.
ODD systems are generally classified into three categories: Retrieval-based Methods, which select responses from a predefined set; Generation-based Methods, which generate responses dynamically using deep learning models; and Hybrid Methods, which combine both retrieval and generation techniques to optimize dialogue quality. Prominent examples of ODD systems include AI-powered chatbots such as OpenAI’s ChatGPT and Google’s Bard, which leverage LLMs to generate contextually rich and dynamic responses. Table 3 summarizes recent advancements in these approaches within the ODD domain.
@@
Self-RAG (asai2023self, ) Self-reflective retrieval guidance with LLMs
@@
ReProver (yang2024leandojo, ) Generation with retrieval-augmented prompting
@@
EMDR2 (singh2021end, ) T5 integration with Top-k MIPS retrieval
@@
LeanDojo (yang2024leandojo, ) Retrieval-augmented theorem proving with LLMs
Chain-of-Note (yu2023chain, ) Enhanced robustness through sequential note taking
@@ 4.1. Retrieval-based Methods
Figure 4. LLMs excel in Open-Domain Dialogue (ODD) by providing accurate answers, understanding emotions, and offering personalized recommendations, making conversations more engaging and tailored to user needs.
In ODD systems, Bordes et al. (bordes2016learning, ) introduced end-to-end learning approaches for dialogue systems, providing an alternative to domain-specific designs. Their work demonstrated methods for learning context-response relationships directly from dialogue data. Henderson et al. (henderson2017efficient, ) implemented these approaches in large-scale commercial systems, while examining scalability requirements and performance characteristics.
Retrieval-based dialogue systems typically operate by selecting appropriate responses from a predefined candidate pool. Given a dialogue context $c$ and a candidate response set $R=\{r_{1},r_{2},...,r_{n}\}$, these systems aim to identify the most suitable response $r^{*}$ according to:
(18) $r^{\ast}=\underset{r_{i}\in R}{\text{argmax}}f(c,r_{i}),$
where $f(c,r_{i})$ represents a matching function that measures the relevance between the context and candidate response.
Early retrieval methods relied primarily on traditional information retrieval techniques such as TF-IDF and BM25. These approaches, while computationally efficient, often struggled to capture semantic relationships and contextual dependencies in dialogue sequences. To address these limitations, researchers began exploring neural retrieval architectures that could better model the semantic space of dialogues.
Karpukhin et al. (karpukhin2020dense, ) presented the Dense Passage Retriever (DPR), which utilizes dense vector representations instead of sparse retrieval methods like TF-IDF and BM25. DPR consists of a dual-encoder framework based on BERT, with separate encoders processing queries and passages to generate dense embeddings. The model training optimizes these embeddings through the following contrastive learning objective:
@@
where $q_{i}$ denotes the query vector, $p_{i}^{+}$ represents the positive passage vector, and $p_{i,j}^{-}$ indicates negative passage vectors. The similarity function $\text{sim}(q,p)$ computes the dot product between embeddings.
For multi-turn dialogue modeling, Tao et al. (tao2019one, ) proposed IoI Network. IoI implements multiple interaction layers between utterances and responses to model semantic relationships. The architecture processes utterance-response pairs sequentially through interaction layers with self-attention mechanisms. The model includes an aggregation component that combines matching signals across layers for response selection. This hierarchical interaction structure enables the model to capture both local and global semantic dependencies within the dialogue context.
Following these developments, several researchers explored methods to enhance the efficiency and effectiveness of neural retrieval models. Yang et al. (yang2024leandojo, ) developed a retrieval-augmented language model framework for contextual information selection. Their method combines dense retrieval with large language models and introduces techniques for identifying relevant context and hard negative examples. The framework includes mechanisms for reducing the search space of candidate premises and implements specialized negative sampling strategies. For multi-turn dialogue modeling, Tao et al. (tao2019one, ) proposed the Interaction-over-Interaction (IoI) Network. IoI implements multiple interaction layers between utterances and responses to model semantic relationships. The architecture processes utterance-response pairs sequentially through interaction layers with self-attention mechanisms. The model includes an aggregation component that combines matching signals across layers for response selection.
These retrieval-based methods operate by leveraging dialogue corpora without requiring explicit annotation of dialogue states or intents. The methods maintain response quality through selection from existing responses, and their modular architecture enables separate optimization of retrieval and ranking components.
Current technical challenges in retrieval-based systems include context dependency modeling, response diversity, and retrieval efficiency at scale. Research directions involve integration of dense retrieval with neural architectures, optimization of negative sampling strategies, and development of efficient retrieval mechanisms for large-scale systems.
@@ 4.2. Generation-based Methods
Generation-based methods in ODD systems aim to produce responses by generating new text rather than selecting from existing responses. These methods have undergone several development stages, from basic sequence modeling to advanced neural architectures. Vinyals and Le (vinyals2015neural, ) and Sutskever et al. (sutskever2014sequence, ) first established sequence-to-sequence frameworks for dialogue generation, demonstrating the feasibility of end-to-end trainable models. Shang et al. (shang2015neural, ) advanced this approach by incorporating attention mechanisms, enabling models to focus on relevant parts of input context during generation. Serban et al. (serban2016building, ) further developed the Hierarchical Recurrent Encoder-Decoder architecture, specifically designed to capture both utterance-level and dialogue-level patterns in conversations. Radford et al. (radford2019language, ) later implemented transformer-based models like GPT-2, introducing self-attention mechanisms and larger-scale pre-training to dialogue generation.
@@ 4.2.1. Knowledge-Enhanced Generation
Knowledge enhancement in dialogue generation addresses the limitation of relying solely on training data by incorporating external information sources. Zhao et al. (zhao2020knowledge, ) combine PLMs with external knowledge through a BERT-based Knowledge Selection Module. This module implements a two-stage process: first selecting relevant documents $D^{\prime}$ from a knowledge base $D$ based on dialogue context $U$, then incorporating this knowledge into response generation. Their GPT-2-based Response Generation Model operates according to:
@@
where $g(U,D^{\prime})$ represents the integration function for dialogue context and selected knowledge, and $\theta$ denotes the GPT-2 parameters. The model implements dynamic knowledge selection during generation, allowing it to adapt knowledge incorporation based on dialogue context.
Asai et al. (asai2023self, ) introduced Self-RAG, a framework incorporating self-reflection mechanisms for knowledge retrieval and response generation. The system implements three key components: a retrieval mechanism that determines when additional knowledge is needed, an evaluation module that assesses retrieved information relevance, and a generation module that produces responses while critiquing its own outputs. Their innovation lies in the use of reflection tokens, which serve multiple purposes: guiding the generation process, enabling runtime behavior customization, and implementing adaptive retrieval thresholds. The framework combines dense retrieval techniques with large language models and introduces specialized negative sampling strategies for improved knowledge selection accuracy. Their approach also includes mechanisms for balancing retrieval frequency and response quality through controlled token generation.
Xu et al. (xu2022long, ) presented PLATO-LTM, featuring a Long-Term Memory (LTM) mechanism for persona information management. The architecture consists of multiple specialized components: a Persona Extractor (PE) based on ERNIE-CNN architecture that processes user input $U_{i}$ to identify persona characteristics, and an LTM module that maintains and retrieves relevant persona information during conversations. Their Generation Module implements a structured learning objective:
(21) $L_{\text{NLL}}=-\mathbb{E}\left[\sum_{t=1}^{T}\log p(r_{t}|c,\rho_{u},\rho_{s},r_{<t})\right],$
where $r_{t}$ denotes the generated response at time $t$, $c$ represents the current context, $\rho_{u}$ and $\rho_{s}$ are user and system persona embeddings respectively, and $r_{<t}$ indicates the response history. The model maintains persona consistency through continuous updates to the LTM component based on interaction patterns.
Song et al. (song2020generate, ) developed the Generate-Delete-Rewrite (GDR) framework for ensuring persona consistency in dialogue generation. The framework implements a three-stage process: initial response prototype generation, identification and masking of persona-inconsistent elements, and output refinement through attention-weighted editing. The prototype generation phase uses a standard encoder-decoder architecture, while the inconsistency detection employs a specialized attention mechanism trained on persona-annotated dialogue data. The refinement stage implements a focused rewriting process that preserves dialogue coherence while enforcing persona constraints.
Tang et al. (tang2023enhancing, ) introduced the Contrastive Latent Variable (CLV) model, advancing personalization through dual processing of sparse and dense persona information. The model first encodes both explicit persona descriptions and implicit persona indicators from dialogue history, then generates responses using a combination of these encoded features. CLV implements a novel clustering mechanism for dense persona descriptions, creating sparse categorical representations that capture essential persona characteristics while reducing computational complexity.